### v1

# The Entropy Flow of a Laser Beam

Abstract
A laser beam is often modelled by a pure coherent state. In fact its state is mixed, even if it has coherent-state photon-number statistics (Poissonian), because the phase must vary. We consider such an ideal laser beam, with phase diffusion rate $\ell$, equal to its (Lorentzian) spectral width. We show that the beam entropy is extensive, with an entropy flow of $\dot{S}=\sqrt{\dot{N}\ell}$, where $\dot{N}$ is the number flow. We give an intuitive explanation for this remarkably simple result, and compare it to a unidirectional thermal beam.
keywords:
entropy flow , laser physics , quantum optics , Feynman-Kac , coherence , phase diffusion ,
journal: Results in Physics
[…]
organization=Quantum and Advanced Technologies Institute, Griffith University,addressline=Yuggera Country, city=Brisbane, postcode=4111, state=Queensland, country=Australia
1 Introduction
Since the invention of the laser, its output has been recognized for, and found myriad applications due to, its superior coherence properties compared to other light sources [Wis16a]. For many purposes, a laser beam may be described as a perfectly coherent state $|{\alpha}\rangle$, with $\alpha$ constant in time, and in some quantum optics textbooks it is so described. However, more careful treatments show that even a laser with with no technical noise has quantum noise leading to phase diffusion, at some rate $\ell$, which is also equal to the FWHMH (full-width-at-half-maximum-height) spread in frequencies [WallsMilb94, MilonniEberly2010]. Since a laser beam is thus not in a pure state, this raises the question: what is its entropy? Surprisingly, to the author’s knowledge, this has never been calculated. This paper rectifies that.
2 Problem statement and solution
A free-space light-beam which is unidirectional, approximately monochromatic, polarized, and fully transverse-coherent is described by a 1-parameter family of bosonic lowering operators $\hat{b}(t)$ [WallsMilb94, BSBW20]. Here, $t$ can be understood as the time since that infinitesimal beam-segment was emitted from the source, so that, at clock-time $\tau$, the field a distance $z$ from the laser is described by $\hat{b}(\tau-z/c)$. With commutation relations $\hat{b}(t),\hat{b}(s)^{\dagger}]=\delta(t-s)$, the operator $\hat{b}^{\dagger}(t)\hat{b}(t)$ is the photon flow operator. We denote the mean photon flow, assumed time-independent, as $\dot{N}=\left\langle{\hat{b}^{\dagger}(t)\hat{b}(t)}\right\rangle$ (rather than $\cal N$} as in Ref. [BSBW20]).
An ideal laser beam can be modelled as a constant-amplitude but wandering-phase coherent state. That is, an eigenstate of $\hat{b}(t)$, with eigenvalue $\sqrt{\dot{N}}e^{i\phi(t)}$, where $\phi(t)$ is a real stochastic process. Separating the beam into small segments of length $\delta t=T/K$, we can write this state over a duration $[0,T)$ as
$\rho_{[0,T)}=\lim_{K\to\infty}\mathbb{E}_{\phi}\left[\bigotimes_{k=0}^{K-1}\left|{re^{i\phi(k\delta t)}}\right\rangle\left\langle{re^{i\phi(k\delta t)}}\right|\right],$ (1)
where $\mathbb{E}_{\bullet}$ denotes the ensemble average over $\bullet$, and the ket $|{re^{i\varphi}}\rangle$ is a single-mode coherent state of mean photon number $r^{2}=\dot{N}\delta t$. We will consider pure phase diffusion,
$\phi(t)=\omega_{0}t+\sqrt{\ell}W(t),$ (2)
where $\omega_{0}\gg\ell$ is the centre-frequency and $W(t)$ is a Wiener process. This well describes quantum-limited phase noise, or other homogenous noise giving rise to a Lorentzian lineshape of FWHMH $\ell$ [MilonniEberly2010]. We also restrict to the usual regime of laser beams where the beam coherence $\mathfrak{C}=4\dot{N}/\ell$ — the number of photons per coherence time [BSBW20] — is extremely large.
We wish to calculate the entropy flow of the beam,
$\dot{S}=\lim_{T\to\infty}T^{-1}S[\rho_{[0,T)}],$ (3)
which is independent of $\phi(0)$ and $T$. Because the von Neumann entropy, $S_{1}$, is too difficult to calculate, we use the Rényi 2-entropy, $S_{2}$. For the asymptotic-in-$T$ results, the beam effectively comprises many statistically independent identical segments of length $\approx 1/\ell$, so these entropies will differ at most by a constant factor [bobkov2018]. Thus, the quantity to be calculated is
$\dot{S}_{2}=\lim_{T\to\infty}T^{-1}\left\{{-k_{\rm B}\log{\rm Tr}\left[(\rho_{[0,T)})^{2}\right]}\right\}$ (4)
As we prove below, Eq. (4), the entropy flow of a laser beam of linewidth $\ell$ much less than its photon flow $\dot{N}$, equals
$\dot{S}_{2}=k_{\rm B}\sqrt{\dot{N}\ell}.$ (5)
3 Proof
Being careful to use independent Wiener processes, say $W(t)$ and $V(t)$, for the two copies of $\rho$ in Eq. (4), we have
${\rm Tr}\left[(\rho_{[0,T)})^{2}\right]=\lim_{K\to\infty}\mathbb{E}_{W,V}\left[{\prod_{k=0}^{K-1}\left|{\left\langle{re^{i\sqrt{\ell}W(k\delta t)}}\right|\left.{re^{i\sqrt{\ell}V(k\delta t)}}\right\rangle}\right|^{2}}\right].$
Evaluating the inner product of these coherent states [WallsMilb94], and defining a new Wiener process $X=(W-V)/\sqrt{2}$, we find
${\rm Tr}\left[(\rho_{[0,T)})^{2}\right]=\lim_{K\to\infty}\mathbb{E}_{X}\left[{\Pi_{k=0}^{K-1}e^{-4r^{2}\sin^{2}\left\{{\sqrt{\ell/2}\,X(k\delta t)}\right\}}}\right].$ (6)
Using $r^{2}=\dot{N}\delta t$, and taking the limit so $\delta t\to{\rm d}t$, gives
${\rm Tr}\left[(\rho_{[0,T)})^{2}\right]=\mathbb{E}_{X}\left[{e^{-4\dot{N}\int_{0}^{T}\sin^{2}\left\{{\sqrt{\ell/2}\,X(t)}\right\}{\rm d}t}}\right].$ (7)
This can be attacked by the Feynman–Kac technique [Kac1949]. Let $X(0)=x$ and define the right-hand-side of Eq. (7) as $u(x,T)$. Then it can be shown [borodin2002handbook, Lalley] that this function is the unique solution, satisfying $u(x,0)=1$, of
[…]
Now, to find $\lambda_{0}$, define a small parameter $\epsilon=(4\mathfrak{C})^{-1/4}=\mbox{$\frac{1}{2}$}(\ell/\dot{N})^{1/4}$ and change variables to $y=\epsilon x$. Then
${\cal L}=2\sqrt{\dot{N}\ell}\left[-\frac{1}{2}\frac{\partial^{2}}{\partial y^{2}}+\frac{1}{2\epsilon^{2}}\sin^{2}(\epsilon y)\right].$ (9)
In the desired high coherence limit, $\mathfrak{C}\gg 1$, the term in square brackets can be approximated by the dimensionless harmonic oscillator Hamiltonian, with minimum eigenvalue of $1/2$. Thus $\lambda_{0}=\sqrt{\dot{N}\ell}$, and we obtain Eq. (5).
4 Discussion of Eq. (5)
First, while such an expression has never been given before, in hindsight it is quite reasonable. The only parameters that define the ideal laser beam are three rates: photon flow $\dot{N}$, linewidth $\ell$, and mean frequency $\omega_{0}$. The entropy of the mixture of coherent states is independent of $\omega_{0}$. The entropy flow must go to zero as $\dot{N}\to 0$ (zero photon flow means a vacuum state, which is pure) and as $\ell\to 0$ (zero linewidth means a constant phase coherent state, which is pure). Therefore $\dot{S}_{2}=\sqrt{\dot{N}\ell}$ is the simplest expression we could hope for.
Second, we can understand Eq. (5) by thinking of the number of distinguishable states in the ensemble (1). Each random walk of the phase through time gives a different state of the beam, but these are not necessarily distinguishable. They are if the coherent states $|{re^{i\phi(k\delta t)}}\rangle$, in some segment $k$, are close to orthogonal. For this we must consider $K$ finite so that the modulus of the coherent amplitude $r=\sqrt{\dot{N}\delta t}$ is non-infinitesimal. Let $\delta t=1/\sqrt{\dot{N}\ell}$, so that $r=(\dot{N}/\ell)^{1/4}$. Over this time, the random change in the phase is, roughly, $\pm\sqrt{\ell\delta t}=\pm(\dot{N}/\ell)^{-1/4}$. Since this is small, the complex coherent amplitude is displaced randomly (on the ring of radius $r$) by $\pm r\sqrt{\ell\delta t}=\pm 1$. This is just the displacement needed to make two coherent states roughly orthogonal. So, every time $\delta t$, a phase-path diverges into two distinguishable phase-paths. This increases the entropy by roughly $k_{\rm B}\log(2)$, giving $\dot{S}\approx k_{\rm B}/\delta t$, in agreement with Eq. (5).
Third, for a ideal standard quantum-noise-limited laser [Wis99], $\ell=\kappa/2\mu$ and $\dot{N}=\kappa\mu$, where $\mu$ is the mean number of intracavity photons and $\kappa$ is the bare cavity intensity-damping rate. This gives $\dot{S}_{2}=\kappa/\sqrt{2}$, clearly limited by quantum noise.
Fourth, it is useful to rewrite Eq. (5) as
$\dot{S}_{2}=k_{\rm B}\sqrt{{P}/{\hbar}}\sqrt{{\ell}/{\omega_{0}}},$ (10)
where $P=\hbar\omega_{0}\dot{N}$ is the laser beam power, to compare with a thermal beam. A unidirectional polarised one-dimensional thermal beam at temperature $\Theta$ has a power [Wis16a]
$P^{\Theta}=\dot{U}^{\Theta}=\int_{0}^{\infty}\frac{{\rm d}\omega}{2\pi}\,\hbar\omega\,\bar{n}^{\Theta}_{\omega}=\frac{\pi}{12}\frac{(k_{\rm B}\Theta)^{2}}{\hbar},$ (11)
where $\bar{n}^{\Theta}_{\omega}=(e^{\hbar\omega/k_{\rm B}\Theta}-1)^{-1}$. Its Rényi-2 entropy flow is
$\dot{S}_{2}^{\Theta}=k_{\rm B}\int_{0}^{\infty}\frac{{\rm d}\omega}{2\pi}\,\log(2\bar{n}^{\Theta}_{\omega}+1)=k_{\rm B}\frac{\pi}{8}\frac{k_{\rm B}\Theta}{\hbar},$ (12)
since a single-mode thermal state has ${\rm Tr}[\rho^{2}]=1/(2\bar{n}+1)$. Equation (12) is just a factor of $3/4$ smaller than the von Neumann entropy flow, $\dot{S}^{\Theta}_{1}$, well known from 1+1–conformal field theory [Calabrese_2009]. Thus, combining Eqs. (11) and (12), we have
$\dot{S}_{2}^{\Theta}=k_{\rm B}\sqrt{{P^{\Theta}}/{\hbar}}\sqrt{{3\pi}/{16}}.$ (13)
Comparing to Eq. (10), we see that, for the same power, the laser beam has an entropy smaller by a factor of roughly $Q^{-1/2}$, where $Q=\omega_{0}/\ell$ is the Q-factor of the beam. Many commercial solid-state lasers have $Q\approx 10^{10}$ or better [Zhao2021]. To give a numerical value, $\dot{S}_{2}\approx 10^{-12}$W/K for $Q=10^{10}$ and P=500P=500mW.
5 Summary and conclusions
Calculating the entropy flow for a laser beam fills a gap in knowledge. The result is remarkably simple, and has an intuitive explanation. It gives a new insight into the nature of laser light and confirms, in a new form, its important differences from (unidirectional) thermal light [Wis16a]. Since a laser is like the pendulum for optical clocks [PhysRevFocus:12.114], the result may also shed light on the question of irreversibility in clocks [Milburn_2020, Oxford2025].
Acknowledgements
This work was supported by Australian Research Council Discovery Project DP220101602. The author thanks Lucas Ostrowski for valuable comments, and acknowledges the use of M365 Copilot to help solve maths problems and find relevant literature. Any results not available from standard mathematical tables of integrals have been checked by the author using theorems or results in the references cited.

### v2

@@
 We consider such an ideal laser beam, with phase diffusion rate $\ell$, equal to its (Lorentzian) spectral width.
-We show that the beam entropy is extensive, with an entropy flow of $\dot{S}=\sqrt{\dot{N}\ell}$, where $\dot{N}$ is the number flow.
-We give an intuitive explanation for this remarkably simple result, and compare it to a unidirectional thermal beam.
+We show that the beam entropy is extensive, with an entropy flow of $\dot{S}=k_{\rm B}\sqrt{\dot{N}\ell}$, where $\dot{N}$ is the number flow.
+We give an intuitive explanation for this remarkably simple result, and compare it to a unidirectional thermal beam’s.
 keywords:
@@
 […]
-organization=Quantum and Advanced Technologies Institute, Griffith University,addressline=Yuggera Country, city=Brisbane, postcode=4111, state=Queensland, country=Australia
+organization=Quantum and Advanced Technologies Research Institute, Griffith University,addressline=Yuggera Country, city=Brisbane, postcode=4111, state=Queensland, country=Australia
 1 Introduction
-Since the invention of the laser, its output has been recognized for, and found myriad applications due to, its superior coherence properties compared to other light sources [Wis16a].
+Since the invention of the laser, its output has been recognized for, and found myriad applications due to, its superior coherence properties compared to other light sources [1].
 For many purposes, a laser beam may be described as a perfectly coherent state $|{\alpha}\rangle$, with $\alpha$ constant in time, and in some quantum optics textbooks it is so described.
-However, more careful treatments show that even a laser with with no technical noise has quantum noise leading to phase diffusion, at some rate $\ell$, which is also equal to the FWHMH (full-width-at-half-maximum-height) spread in frequencies [WallsMilb94, MilonniEberly2010].
+However, more careful treatments show that even a laser with with no technical noise has quantum noise leading to phase diffusion, at some rate $\ell$, which is also equal to the FWHMH (full-width-at-half-maximum-height) spread in frequencies [2, 3].
 Since a laser beam is thus not in a pure state, this raises the question: what is its entropy?
@@
 2 Problem statement and solution
-A free-space light-beam which is unidirectional, approximately monochromatic, polarized, and fully transverse-coherent is described by a 1-parameter family of bosonic lowering operators $\hat{b}(t)$ [WallsMilb94, BSBW20].
+A free-space light-beam which is unidirectional, approximately monochromatic, polarized, and fully transverse-coherent is described by a 1-parameter family of bosonic lowering operators $\hat{b}(t)$ [2, 4].
 Here, $t$ can be understood as the time since that infinitesimal beam-segment was emitted from the source, so that, at clock-time $\tau$, the field a distance $z$ from the laser is described by $\hat{b}(\tau-z/c)$.
@@
 We denote the mean photon flow, assumed time-independent, as $\dot{N}=\left\langle{\hat{b}^{\dagger}(t)\hat{b}(t)}\right\rangle$ (rather than $\cal N$} as in Ref.
-[BSBW20]).
+[4]).
 An ideal laser beam can be modelled as a constant-amplitude but wandering-phase coherent state.
@@
 We will consider pure phase diffusion,
-$\phi(t)=\omega_{0}t+\sqrt{\ell}W(t),$ (2)
-where $\omega_{0}\gg\ell$ is the centre-frequency and $W(t)$ is a Wiener process.
-This well describes quantum-limited phase noise, or other homogenous noise giving rise to a Lorentzian lineshape of FWHMH $\ell$ [MilonniEberly2010].
-We also restrict to the usual regime of laser beams where the beam coherence $\mathfrak{C}=4\dot{N}/\ell$ — the number of photons per coherence time [BSBW20] — is extremely large.
+$\phi(t)=-\omega_{0}t+\sqrt{\ell}W(t),$ (2)
+where $\omega_{0}\ (\gg\ell)$ is the centre-frequency and $W(t)$ is a Wiener process.
+This well describes quantum-limited phase noise, or other homogenous noise giving rise to a Lorentzian lineshape of FWHMH $\ell$ [3].
+We also restrict to the usual regime of laser beams where the beam coherence $\mathfrak{C}=4\dot{N}/\ell$ — the number of photons per coherence time [4] — is extremely large.
 We wish to calculate the entropy flow of the beam,
@@
 Because the von Neumann entropy, $S_{1}$, is too difficult to calculate, we use the Rényi 2-entropy, $S_{2}$.
-For the asymptotic-in-$T$ results, the beam effectively comprises many statistically independent identical segments of length $\approx 1/\ell$, so these entropies will differ at most by a constant factor [bobkov2018].
+For the asymptotic-in-$T$ results, the beam effectively comprises many statistically independent identical segments of length $\approx 1/\ell$, so these entropies will differ at most by a constant factor [5].
 Thus, the quantity to be calculated is
@@
 ${\rm Tr}\left[(\rho_{[0,T)})^{2}\right]=\lim_{K\to\infty}\mathbb{E}_{W,V}\left[{\prod_{k=0}^{K-1}\left|{\left\langle{re^{i\sqrt{\ell}W(k\delta t)}}\right|\left.{re^{i\sqrt{\ell}V(k\delta t)}}\right\rangle}\right|^{2}}\right].$
-Evaluating the inner product of these coherent states [WallsMilb94], and defining a new Wiener process $X=(W-V)/\sqrt{2}$, we find
+Evaluating the inner product of these coherent states [2], and defining a new Wiener process $X=(W-V)/\sqrt{2}$, we find
 ${\rm Tr}\left[(\rho_{[0,T)})^{2}\right]=\lim_{K\to\infty}\mathbb{E}_{X}\left[{\Pi_{k=0}^{K-1}e^{-4r^{2}\sin^{2}\left\{{\sqrt{\ell/2}\,X(k\delta t)}\right\}}}\right].$ (6)
@@
 ${\rm Tr}\left[(\rho_{[0,T)})^{2}\right]=\mathbb{E}_{X}\left[{e^{-4\dot{N}\int_{0}^{T}\sin^{2}\left\{{\sqrt{\ell/2}\,X(t)}\right\}{\rm d}t}}\right].$ (7)
-This can be attacked by the Feynman–Kac technique [Kac1949].
+This can be attacked by the Feynman–Kac technique [6].
 Let $X(0)=x$ and define the right-hand-side of Eq.
 (7) as $u(x,T)$.
-Then it can be shown [borodin2002handbook, Lalley] that this function is the unique solution, satisfying $u(x,0)=1$, of
+Then it can be shown [7, 8] that this function is the unique solution, satisfying $u(x,0)=1$, of
 […]
 […]
-Now, to find $\lambda_{0}$, define a small parameter $\epsilon=(4\mathfrak{C})^{-1/4}=\mbox{$\frac{1}{2}$}(\ell/\dot{N})^{1/4}$ and change variables to $y=\epsilon x$. Then
+Now, to find $\lambda_{0}$, define a small parameter $\epsilon=(4\mathfrak{C})^{-1/4}=\mbox{$\frac{1}{2}$}(\ell/\dot{N})^{1/4}$ and change variables to $y=\sqrt{\ell/2}\,x/\epsilon$. Then
 ${\cal L}=2\sqrt{\dot{N}\ell}\left[-\frac{1}{2}\frac{\partial^{2}}{\partial y^{2}}+\frac{1}{2\epsilon^{2}}\sin^{2}(\epsilon y)\right].$ (9)
@@
 The entropy flow must go to zero as $\dot{N}\to 0$ (zero photon flow means a vacuum state, which is pure) and as $\ell\to 0$ (zero linewidth means a constant phase coherent state, which is pure).
-Therefore $\dot{S}_{2}=\sqrt{\dot{N}\ell}$ is the simplest expression we could hope for.
+Therefore $\dot{S}_{2}=k_{\rm B}\sqrt{\dot{N}\ell}$ is the simplest expression we could hope for.
 Second, we can understand Eq.
@@
 (5).
-Third, for a ideal standard quantum-noise-limited laser [Wis99], $\ell=\kappa/2\mu$ and $\dot{N}=\kappa\mu$, where $\mu$ is the mean number of intracavity photons and $\kappa$ is the bare cavity intensity-damping rate.
-This gives $\dot{S}_{2}=\kappa/\sqrt{2}$, clearly limited by quantum noise.
+Third, for a ideal standard quantum-noise-limited laser [9], $\ell=\kappa/2\mu$ and $\dot{N}=\kappa\mu$, where $\mu$ is the mean number of intracavity photons and $\kappa$ is the bare cavity intensity-damping rate.
+This gives $\dot{S}_{2}/k_{\rm B}=\kappa/\sqrt{2}$, clearly limited by quantum noise.
 Fourth, it is useful to rewrite Eq. (5) as
 $\dot{S}_{2}=k_{\rm B}\sqrt{{P}/{\hbar}}\sqrt{{\ell}/{\omega_{0}}},$ (10)
-where $P=\hbar\omega_{0}\dot{N}$ is the laser beam power, to compare with a thermal beam. A unidirectional polarised one-dimensional thermal beam at temperature $\Theta$ has a power [Wis16a]
+where $P=\hbar\omega_{0}\dot{N}$ is the laser beam power, to compare with a thermal beam. A unidirectional polarised one-dimensional thermal beam at temperature $\Theta$ has a power [1]
 $P^{\Theta}=\dot{U}^{\Theta}=\int_{0}^{\infty}\frac{{\rm d}\omega}{2\pi}\,\hbar\omega\,\bar{n}^{\Theta}_{\omega}=\frac{\pi}{12}\frac{(k_{\rm B}\Theta)^{2}}{\hbar},$ (11)
@@
 since a single-mode thermal state has ${\rm Tr}[\rho^{2}]=1/(2\bar{n}+1)$.
-Equation (12) is just a factor of $3/4$ smaller than the von Neumann entropy flow, $\dot{S}^{\Theta}_{1}$, well known from 1+1–conformal field theory [Calabrese_2009].
+Equation (12) is just a factor of $3/4$ smaller than the von Neumann entropy flow, $\dot{S}^{\Theta}_{1}$, well known from 1+1–conformal field theory [10].
 Thus, combining Eqs.
@@
 (10), we see that, for the same power, the laser beam has an entropy smaller by a factor of roughly $Q^{-1/2}$, where $Q=\omega_{0}/\ell$ is the Q-factor of the beam.
-Many commercial solid-state lasers have $Q\approx 10^{10}$ or better [Zhao2021].
-To give a numerical value, $\dot{S}_{2}\approx 10^{-12}$W/K for $Q=10^{10}$ and P=500P=500mW.
+Many commercial solid-state lasers have $Q\approx 10^{10}$ or better [11].
+To give a numerical value, for $Q=10^{10}$ and P=750P=750mW we have $\dot{S}_{2}\approx 10^{-12}$W/K or, perhaps more usefully, $\sqrt{\dot{N}\ell}\approx 10^{11}{\rm s}^{-1}$.
 5 Summary and conclusions
@@
 The result is remarkably simple, and has an intuitive explanation.
-It gives a new insight into the nature of laser light and confirms, in a new form, its important differences from (unidirectional) thermal light [Wis16a].
-Since a laser is like the pendulum for optical clocks [PhysRevFocus:12.114], the result may also shed light on the question of irreversibility in clocks [Milburn_2020, Oxford2025].
+It gives new insight into the nature of laser light and confirms, in a new form, its important differences from (unidirectional) thermal light [1].
+Since a laser is like the pendulum for optical clocks [12], the result may also shed light on the question of irreversibility in clocks [13, 14].
 Acknowledgements
 This work was supported by Australian Research Council Discovery Project DP220101602.
-The author thanks Lucas Ostrowski for valuable comments, and acknowledges the use of M365 Copilot to help solve maths problems and find relevant literature.
+The author thanks Lucas Ostrowski and Kiarn Laverick for valuable comments, and acknowledges the use of M365 Copilot to help solve maths problems and find relevant literature.
 Any results not available from standard mathematical tables of integrals have been checked by the author using theorems or results in the references cited.